# EVE: Towards End-to-End Video Subtitle Extraction with Vision-Language Models

### v1

@@ Abstract
The advent of Large Vision-Language Models (LVLMs) has advanced the video-based tasks, such as video captioning and video understanding. Some previous research indicates that taking texts in videos as input can further improve the performance of video understanding. As a type of indispensable information in short videos or movies, subtitles can assist LVLMs to better understand videos. Most existing methods for video subtitle extraction are based on a multi-stage framework, handling each frame independently. They can hardly exploit the temporal information of videos. Although some LVLMs exhibit the robust OCR capability, predicting accurate timestamps for subtitle texts is still challenging. In this paper, we propose an End-to-end Video Subtitle Extraction method, called EVE, which consists of three modules: a vision encoder, an adapter module, and a large language model. To effectively compress the visual tokens from the vision encoder, we propose a novel adapter InterleavedVT to interleave two modalities. It contains a visual compressor and a textual region compressor. The proposed InterleavedVT exploits both the merits of average pooling and Q-Former in token compression. Taking the temporal information of videos into account, we introduce a sliding-window mechanism in the textual region compressor. To benchmark the video subtitle extraction task, we propose a large dataset ViSa including 2.5M videos. Extensive experiments on ViSa demonstrate that the proposed EVE can outperform existing open-sourced tools and LVLMs.
1 Introduction
Due to the current prevalence of video as a mainstream media format, a significant amount of research is centered around video-related studies and developments, such as video understanding [34, 35, 37]. In the video understanding task, the text information present in videos, especially subtitle, plays an indispensable role since it provides more contextually relevant information of videos. Accurately extracting subtitles from videos can significantly assist existing Vision-Language Models (VLMs) in better understanding video content. Furthermore, video subtitle extraction can also provide foundational support for practical applications such as movie subtitle translation.
@@ 1 Introduction
Figure 1: Most previous methods (a) rely on a multi-stage framework, where prediction error accumulates at each stage, influencing the final prediction. In contrast, the proposed EVE (b) is an end-to-end video subtitle extraction method.
@@
Previous video subtitle extraction methods can be broadly categorized into two groups: OCR-based approach and speech-based approach. The OCR-based approach [7] often involves multiple stages, such as text detection, subtitle classification, and text recognition for subtitle regions (as shown in Fig. 1 (a)). Although these methods achieve reasonable performance, the multi-stage approach is susceptible to error accumulation. For example, if the output of text detection is inaccurate, text recognition models can hardly recognize corresponding texts correctly. In addition, they process each frame independently without considering the temporal information between frames. Different from the OCR-based approach, the speech-based approach [23, 22] directly transcribes the speech of videos to texts and timestamps the texts to generate subtitles. However, these methods are sensitive to noise in audios and often struggle to align the recognized subtitles with visual text cues in videos.
Recent advancements in VLMs have shown improved performance across various video tasks [12, 32, 38, 36]. Nevertheless, the field of video subtitle extraction still lacks an efficient end-to-end solution and corresponding datasets. In this paper, we introduce an End-to-end Video Subtitle Extraction method EVE, consisting of a vision encoder, an adapter module, and a Large Language Model (LLM). Specifically, we employ Swin-Transformer [18] as the vision encoder to extract visual features of each frame. Subsequently, we develop an innovative adapter module, named InterleavedVT, to compress the output visual tokens of the vision encoder. The proposed InterleavedVT consists of a visual compressor and a textual region compressor. The visual compressor is based on an average pooling operation to provide the global positional and visual information of subtitles within each video frame while the textual region compressor leverages a Q-Former [14] approach to extract subtitle-related tokens. Then, the output tokens of these two compressors are interleaved to construct a token sequence, which is further input into an LLM to predict subtitles. To support bilingual video subtitle extraction, we adopt Qwen2 [40] as the language model of our method. Additionally, to benchmark this task, we propose a large Video Subtitle Extraction dataset, called ViSa, based on existing two video datasets CNVid [9] and MovieChat [27]. The proposed ViSa includes 2.35M videos for pre-training, 230K videos for supervised fine-tuning, and 1.6K videos for test. The experimental results on ViSa demonstrate that EVE can achieve a significant performance improvement compared with previous open-sourced methods.
In summary, the contributions of this paper are as follows:
  • We first propose an end-to-end video subtitle extraction model EVE, consisting a vision encoder, an adapter module, and a large language model.
  • To condense the tokens of each frame, we develop a novel adapter InterleavedVT for video subtitle extraction. In InterleavedVT, a visual compressor and a textual region compressor are used to extract the global visual information and subtitle-related text information, respectively.
  • To benchmark the video subtitle extraction task, we propose a large video subtitle extraction dataset, called ViSa, including over 2.5M videos.
  • Compared with previous video subtitle extraction methods, the proposed EVE can achieve better performance in the test set of ViSa.
2 Related Work
2.1 Video Subtitle Extraction
Subtitles, as supplementary information in videos, play a crucial role in video understanding. Existing approaches to video subtitle extraction are mainly divided into two categories: OCR-based approach and speech-based approach. OCR-based methods typically treat video subtitle extraction as a multi-stage task. For instance,  [7] first perform text localization to locate the region of interest of frames. Then, it sequentially detects and recognizes subtitle texts in target frames.  [39] also adopts a similar pipeline, where Connectionist Text Proposal Network (CTPN) [28] is utilized for video subtitle detection. These methods consider information of each frame independently, struggling to leverage contextual information of videos. Additionally, errors in text detection and recognition will impact subtitle prediction. Speech-based methods [24] directly feed the audio content of videos into a speech recognition model, and further incorporate the transcribed texts with timestamps to generate corresponding subtitles. Since these methods treat the transcribed texts of the speech directly as subtitles, leading to discrepancies with the actual subtitles in videos. Especially in short videos (such as YouTube Shorts and TikTok videos), the alignment between subtitles and speech may not be perfect. Moreover, some open-source tools (e.g., VideOCR [^1] https://github.com/devmaxxing/videocr-PaddleOCR and PyVideoTrans [^2] https://github.com/jianchang512/pyvideotrans) or commercial cloud services (e.g., OpenVision [^3] https://vision.aliyun.com/ and GhostCut [^4] https://cn.jollytoday.com/home/) also support video subtitle extraction.
@@ 2.1 Video Subtitle Extraction
Figure 2: Overall architecture of EVE (a) and details of InterleavedVT (b). EVE contains three modules: a vision encoder, an adapter and a large language model. We propose a novel adapter called InterleavedVT, consisting of a visual compressor and a textual region compressor. To enhance the correlations between adjacent frames, we introduce a slide window mechanism.
2.2 Video Token Merging
Different from the tasks that take one single image as input, video-related tasks tend to involve more frames. Therefore, video-related Vision-Language Models (VLMs) need to employ token merging or compressing techniques to increase the number of input frames [13, 31, 45, 42]. Inspired by ToMe [5], many works have proposed various token merging methods to reduce the computational cost of models [16, 17]. Since video-related tasks often require leveraging the pre-trained vision encoder from image tasks, video-related tasks typically avoid compressing token quantities within the vision encoder and instead process the visual tokens output by the vision encoder. A common method is to employ a Q-Former-like module [14] to perform cross-attention operation between $N$ query tokens and visual tokens resulting $N$ compressed tokens, e.g., LLaMA-VID [15], TimeChat [26] and Video-LLaMA [43]. However, a recent research DeCo [41] indicates that only using Q-Former to compress visual tokens requires a large amount of data for training and may not converge well. In contrast, a simple average pooling may even perform better. Thus, some methods directly adopt an average pooling layer to compress input tokens, such as LITA [10], Video-ChatGPT [21] and Oryx MLLM [19]. To combine the metrits of both approaches, we propose Interleave VT, which achieves a high token compression rate while maintaining strong generalizability.
3 Methodology
3.1 Architecture
The overall architecture of the proposed End-to-end Video Subtitle Extraction method EVE is shown in Figure 2. Specifically, EVE consists of three modules: a vision encoder, an adapter and a large language model. Given an input video $\mathcal{V}=\{\mathbf{I}_{0},...,\mathbf{I}_{N-1}\}$, each frame $\mathbf{I}_{i}$ ($i\in\{0,1,...,N-1\}$) is independently input into the vision encoder to extract corresponding visual features $\mathbf{F}_{i}$. Thus, the output of the vision encoder can be denoted as =$\mathcal{F}=\{\mathbf{F}_{0},...,\mathbf{F}_{N-1}\}$, where $\mathbf{F}_{i}\in\mathbb{R}^{(H\times W)\times C}$ denotes the visual tokens of the $i^{th}$ frame. Subsequently, the visual features of all frames $\mathcal{F}$ are input into the proposed novel adapter InterleavedVT to produce the compressed visual tokens ^=$\hat{\mathcal{F}}=\{\hat{\mathbf{F}}_{0},...,\hat{\mathbf{F}}_{N-1}\}$. Finally, the visual tokens $\hat{\mathcal{F}}$ and embedded prompt $\mathcal{P}$ are fed into the large language model to generate corresponding subtitles $\mathcal{S}=\{S_{0},...,S_{T-1}\}$. The template of subtitle $S_{i}$ is “$\langle b_{i}\rangle\langle e_{i}\rangle s_{i}$”, where $b_{i}$/$e_{i}$ represents the begin/end index of frame for the $i$-th subtitle and $s_{i}$ is the corresponding text. In the following, we will detail each module in the proposed EVE.
@@ 3.1 Architecture
Vision Encoder. In recent years, various Transformer-based vision backbones are proposed, including ViT [6], Swin-Transformer [18], EViT [16], Dynamic ViT [25], etc. Among these Transformer-based backbones, we select Swin-Transformer pretrained in Donut [11] as the vision encoder of EVE for two reasons: (1) Better initialization for OCR. Since the Swin-Transformer backbone of Donut has been pre-trained on numerous document datasets for multiple OCR tasks, it can provide better initialization to enhancing the text perception ability of EVE. In addition, based on the pre-training Swin-Transformer of Donut, our model can converge faster during the pre-training phase. (2) Higher time efficiency. Benefiting from the introduced shifted window-based self-attention module in Swin-Transformer, it shows better time efficiency, which can accelerate extracting visual features of frames in videos.
InterleavedVT. Although previous methods have proposed various adapter for video tasks, they may not suit for video subtitle task. To effectively compress the output visual tokens $\mathcal{F}$ in the video subtitle extraction task, we propose an innovative adapter InterleavedVT, which consists of a visual compressor and a textual region compressor. The details of InterleavedVT are shown in Figure 2. InterleavedVT fuses two types of token compression methods: average pooling and Q-Former. Due to the consistency of the subtitle positions in the video, the visual compressor is introduced to perform an average pooling operation on $\mathbf{F}_{i}$ to reserve the global visual features $\mathbf{F}^{v}_{i}\in\mathbb{R}^{4\times 4\times C}$ with relative positions. In addition, considering that recognizing texts of subtitles requires fine-grained textual information, we introduce the textual region compressor to reserve text-related tokens. Specifically, we randomly initialize $K$ queries $\mathbf{Q}$ to performance cross-attention with $\mathbf{F}_{i}$, resulting $K$ text-related tokens $\mathbf{F}^{t}_{i}\in\mathbb{R}^{K\times C}$, which can be defined as:
$\mathbf{F}^{t}_{i}=\texttt{CrossAttention}(Q,\mathbf{F}_{i})$ (1)
Taking into account the repeated subtitles between adjacent frames, we introduced a window mechanism in the textual region compressor to enhance the robustness of text-related token extraction. Thus, the processing of the textual region compressor can be modified into:
$\mathbf{F}^{t}_{i}=\texttt{CrossAttention}(Q,[\mathbf{F}_{i-1},\mathbf{F}_{i},\mathbf{F}_{i+1}])$ (2)
Similar to previous methods [26, 8], a projection layer is needed to align the features of vision encoder and language models. However, we observe that utilizing a single projection layer for both the visual and text-related tokens may lead to subpar performance (more discussions ). Since the two types of features from the visual compressor and the textual region compressor are regarded as different representations of one frame, we employ two projection layers to transform $\mathbf{F}^{v}_{i}$ and $\mathbf{F}^{t}_{i}$ into $\hat{\mathbf{F}}^{v}_{i}$ and $\hat{\mathbf{F}}^{t}_{i}$, respectively. Finally, the global visual tokens ^$\hat{\mathcal{F}}^{v}$ and the text-related tokens ^$\hat{\mathcal{F}}^{t}$ of all frames are interleaved into the output of InterleavedVT $\mathbf{F}^{out}$:
$\mathbf{F}^{out}=[\hat{\mathbf{F}}^{v}_{0},\hat{\mathbf{F}}^{t}_{0},...,\hat{\mathbf{F}}^{v}_{N-1},\hat{\mathbf{F}}^{t}_{N-1}]$ (3)

### v2

@@ Abstract
Video subtitles play a crucial role in short videos and movies, as they not only help models better understand video content but also support applications such as video translation and content retrieval. Existing video subtitle extraction methods typically rely on multi-stage frameworks, where errors accumulate across stages and temporal dependencies are underutilized due to frame-wise processing. Moreover, although some Large Vision-Language Models (LVLMs) possess strong OCR capabilities, predicting accurate timestamps for subtitle texts remains challenging. To this end, we propose an End-to-end Video subtitle Extraction framework based on LVLMs, named EVE, which can output subtitles and their timestamps simultaneously. Specifically, we introduce a dual-branch Spatiotemporal Subtitle-Salient (S3) Module that serves as an adapter for LVLMs, capable of representing subtitle-related content and considering inter-frame correlations using only a small number of tokens. Within this module, the Spatial Semantic Context Aggregate branch aggregates high-level global semantics to provide spatial visual contextual information, while the Temporal Subtitle Token Query branch explicitly queries subtitle-relevant tokens while considering temporal correlation across frames. The small number of tokens retained by the S3 module are fed to the language model, which then directly outputs the subtitle text along with its timestamps. Furthermore, we construct the first large-scale dataset dedicated to video subtitle extraction, ViSa, containing over 2.5M videos with timestamped and bilingual annotation, thereby providing the community with a well-organized training and evaluation benchmark. Extensive experiments on ViSa demonstrate that EVE significantly outperforms existing open-source tools and LVLMs in both subtitle recognition accuracy and temporal alignment, achieving state-of-the-art performance.
I Introduction
With the rapid growth of video as the dominant medium for information dissemination, video understanding has become a central topic in multi-modal research [wu2021towards, wu2024deep, tip-video-sum1, tip-video-sum2]. Among the various modalities embedded in videos, textual information—particularly subtitles—serves as an indispensable bridge between visual and linguistic domains, providing explicit and contextually aligned semantic information. Accurate subtitle extraction not only enhances the ability of Vision-Language Models (VLMs) to comprehend video content but also enables a wide range of downstream applications, such as cross-lingual subtitle translation and video content retrieval [tip-video-ret]. Therefore, video subtitle extraction is of great importance both as a fundamental research problem and as a key enabling technology for practical video understanding applications.
@@ I Introduction
Figure 1: Comparison of the proposed method with previous methods on the pipeline. Most previous methods (a) rely on a multi-stage framework, where prediction error accumulates at each stage, influencing the final prediction. In contrast, the proposed EVE (b) is an end-to-end video subtitle extraction method.
@@
Previous video subtitle extraction methods can be broadly categorized into two groups: OCR-based approach and speech-based approach. The OCR-based approach [elshahaby2022end, yan2020end] often involves multiple stages, such as text detection, temporal tracking, and text recognition for subtitle regions as shown in Figure 1 (a). Although these methods achieve reasonable performance, the multi-stage approach is susceptible to error accumulation. For example, if the output of text detection is inaccurate, text recognition models can hardly recognize the corresponding texts correctly. In addition, they process each frame independently without considering the temporal information between frames. Different from the OCR-based approach, the speech-based approach [papi2023direct, mathur2015generating] directly transcribes the speech of videos to texts and timestamps the texts to generate subtitles. However, these methods are sensitive to noise in audios and often struggle to align the recognized subtitles with visual text cues in videos.
Recent advances in Vision-Language Models (VLMs) have substantially improved the performance of various video-related tasks [kim2024image, wang2024videotree, xu2024pllava, wu2023bidirectional]. However, the task of video subtitle extraction remains underexplored. To our knowledge, there is no efficient end-to-end VLM-based frameworks or comprehensive datasets available. In addition, video subtitle extraction poses two unique challenges that distinguish it from general video understanding. First, subtitles appear in very small spatial regions and depend on fine-grained visual cues, which are easily lost under the token compression and global pooling strategies typical of existing LVLMs. Second, subtitles persist across multiple frames with gradual transitions, requiring high-fidelity modeling of both their textual content and temporal boundaries. Such temporal coherence cannot be captured by frame-wise OCR or captioning, often resulting in inconsistent or misaligned subtitles.
To fill this gap, we propose an End-to-end Video subtitle Extraction framework, termed EVE. EVE comprises three major components: a vision encoder, a Spatiotemporal Subtitle-Salient (S3) module, and a large language model (LLM). The S3 module consists of two branch and acts as an adapter that refines frame-level visual features from two complementary perspectives while retaining only a small number of tokens. Specifically, the Spatial Semantic Context Aggregate (SSCA) branch aggregate high-level global semantics to provide spatial contextual cues for subtitle extraction. Meanwhile, the Temporal Subtitle Token Query (TSTQ) branch explicitly queries subtitle-relevant tokens inspired by the Q-former [li2023blip]. Q-former is designed for general vision-language alignment, while TSTQ explicitly targets subtitle-bearing regions and enforces temporal consistency. The outputs from both branches are individually projected into the language model’s embedding space and then concatenated before being fed into the LLM to generate timestamped subtitle texts. To support bilingual (i.e., Chinese and English) video subtitle extraction, we employ Qwen2 [yang2024qwen2] as the large language model, leveraging its strong cross-lingual understanding and generation capabilities to ensure subtitle accuracy.
Although several datasets [elshahaby2022end, 1227749, 6628859, Wang20111457] have been used in existing video subtitle extraction method, they suffer from issues such as limited scale, lack of timestamp annotations, or derivation from images rather than videos. This renders these datasets vastly inferior to those for other relevant tasks, rendering them inadequate to meet the current data requirements of LVLM. Therefore, we construct a large-scale benchmark dataset, ViSa, containing over 2.5M videos collected from CNVid [gan2023cnvid] and MovieChat [song2024moviechat] for pre-training, supervised fine-tuning, and testing. ViSa provides bilingual annotations of subtitle texts and timestamps, offering a valuable resource for systematic evaluation and future research on video subtitle extraction.
Extensive experiments on ViSa demonstrate that EVE significantly outperforms existing open-source tools and LVLM baselines in both subtitle recognition accuracy and temporal alignment, setting a new state-of-the-art for video subtitle extraction. In summary, the main contributions of this work are as follows:
  • We propose an End-to-end Video subtitle Extraction framework, EVE, built upon LVLMs to directly generate timestamped subtitles from video input.
  • We design a Spatiotemporal Subtitle-Salient (S3) Module, which refines visual features through spatial semantic aggregation and temporal subtitle token querying to effectively capture subtitle-salient representations and consider temporal consistency with only a small number of tokens.
  • We construct the first large-scale dataset, ViSa, containing over 2.5M videos with bilingual subtitle text and timestamp annotations, providing a valuable resource for systematic evaluation and future research in this field.
  • Extensive experiments on ViSa demonstrate that EVE significantly outperforms existing open-source tools and LVLM baselines in video subtitle extraction task.
The remainder of this paper is organized as follows: Section II reviews existing work on video subtitle extraction. Section III presents the overall architecture of EVE, including the vision encoder, the proposed S3 module, and the training and inference procedures. Section IV outlines the characteristics of the ViSa dataset and its construction process. Section V reports extensive experiments, including comparisons with existing tools and LVLMs, ablation studies, visualizations, and hyperparameter analyses. Finally, Section VI concludes the paper, discusses the limitations of the proposed method, and outlines potential future directions.
II Related Work
II-A Video Subtitle Extraction
Subtitles, as supplementary information in videos, play a crucial role in video understanding. Existing approaches to video subtitle extraction can be broadly categorized into two groups: OCR-based and speech-based approaches.
II-A1 OCR-based approach
OCR-based approaches typically treat subtitle extraction as a multi-stage pipeline. For example, Elshahaby et al.  [elshahaby2022end] proposed an system that first performs text localization to identify regions of interest within video frames, followed by text detection and recognition on the localized areas. Similarly, Yan et al. [yan2020end] adopted a comparable framework, employing the Connectionist Text Proposal Network (CTPN) [tian2016detecting] for subtitle region detection. However, such methods process each frame independently, making it difficult to leverage the temporal and contextual information across frames. Moreover, the sequential nature of detection and recognition stages often leads to error accumulation, where inaccurate detection can significantly degrade the final recognition performance. VideOCR [^1] https://github.com/devmaxxing/videocr-PaddleOCR is an open-source video subtitle extraction tool developed based on the OCR tool PaddleOCR  [^2] https://github.com/PaddlePaddle/PaddleOCR.
II-A2 Speech-based approach
Mathur et al. [mathur2015generating] presents an offline pipeline that extracts audio from video, applies speech recognition, and converts the recognized speech into time-aligned subtitle files. Ramani et al.  [ramani2020automatic] extract subtitles by transcribing the audio stream of videos using speech recognition models and then associating the transcribed text with timestamps. Sara et al. [papi2023direct] proposes an end-to-end direct speech translation model that generates time-aligned target-language subtitles from source audio. While effective for clear speech, these methods directly treat transcribed speech as subtitles, which can cause mismatches with the actual on-screen text—particularly in short-form videos (e.g., YouTube Shorts, TikTok) where subtitle alignment with speech is often imperfect.
In addition to the aforementioned two types of methods, there are also tools available for extracting video subtitles. PyVideoTran  [^3] https://github.com/jianchang512/pyvideotrans is a fully automatic video/audio translation tool that transcribes speech, generates and translates subtitles, which also supports batch processing of video or audio files into precisely time-coded SRT subtitle files. Commercial cloud services (e.g., OpenVision [^4] https://vision.aliyun.com/, GhostCut [^5] https://cn.jollytoday.com/home/) have also been developed for video subtitle extraction.
II-B Large Vision-Language Models
In recent years, Large Vision-Language Models (LVLMs) [wang2024qwen2, yao2024minicpm, zhang2023video, li2025llama, huang2025lita] have demonstrated remarkable capabilities in instruction following and multimodal reasoning, enabling them to capture on-screen text or subtitles when prompted with appropriate instructions. Unlike conventional OCR-based methods, these models can leverage global visual-semantic context and learn high-level associations between video content and textual cues. However, their subtitle extraction performance remains limited in both accuracy and temporal alignment, mainly due to the lack of task-specific training for fine-grained subtitle localization. Most existing LVLMs are optimized for holistic video captioning or question answering [song2024moviechat, wang2024qwen2] rather than precise frame-level text prediction. Consequently, they tend to produce scene-level descriptions or paraphrased summaries instead of exact subtitle transcripts. This highlights the need for a task-specific framework that can jointly model spatial and temporal semantics to achieve accurate and temporally consistent subtitle extraction.
@@ II-B Large Vision-Language Models
Figure 2: Overall architecture of EVE (a) and details of S3 module (b). EVE contains three modules: a vision encoder, an adapter and a large language model. We propose a novel adapter called S3 module, consisting of a spatial semantic context aggregating and a temporal subtitle token querying branches. To enhance the correlations between adjacent frames, we also introduce a temporal window mechanism.
II-C Video Token Compression
The proposed S3 module aims to extract subtitle-relevant representations from dense visual tokens and retain only a compact set of key tokens for the large language model. This mechanism is conceptually related to the token compression strategies widely adopted in recent LVLMs, which also focus on preserving the most informative tokens while reducing redundancy to enhance efficiency. Existing token compression methods can be roughly divided into two categories. The first category relies on token merging, where similar visual tokens are grouped or averaged to reduce redundancy, as exemplified by ToMe [bolyatoken] and its variants [liang2022not, liusimple]. The second category employs query-based token selection, which introduces a small set of learnable queries to extract compact visual representations through cross-attention, similar to the Q-Former module [li2023blip]. This approach has been adopted by several video-related VLMs such as LLaMA-VID [li2025llama], TimeChat [ren2024timechat], and Video-LLaMA [zhang2023video]. However, studies such as DeCo [yao2024deco] have shown that Q-Former-based compression often requires large-scale training data and can suffer from unstable convergence. In contrast, simple operations like average pooling have demonstrated surprisingly strong performance in some cases, as reported in LITA [huang2025lita], Video-ChatGPT [maaz2023video], and Oryx  [liu2024oryx]. Although our S3 module shares the same philosophy of extracting a minimal yet informative token set, it is specifically designed for the video subtitle extraction task. Instead of general-purpose visual redundancy reduction, S3 focuses on identifying tokens that are semantically and temporally relevant to subtitles.
III Methodology
III-A Architecture
The overall architecture of the proposed End-to-end Video subtitle Extraction framework (EVE) is illustrated in Figure 2. EVE consists of three main components: a vision encoder, a Spatiotemporal Subtitle-Salient (S3) Module, and a large language model (LLM). Given an input video $\mathcal{V}=\{\mathbf{I}_{0},...,\mathbf{I}_{N-1}\}$, each frame $\mathbf{I}_{i}$ ($i\in\{0,1,...,N-1\}$) is first processed by the vision encoder to extract frame-level visual features $\mathbf{F}_{i}$. The resulting feature sequence can be denoted as $\mathcal{F}=\{\mathbf{F}_{0},...,\mathbf{F}_{N-1}\}$, where $\mathbf{F}_{i}\in\mathbb{R}^{(H\times W)\times C}$ represents the dense visual tokens of the $i$-th frame. To obtain subtitle-relevant representations, we propose the S3 module, which comprises two components: the Spatial Semantic Context Aggregation (SSCA) and the Temporal Subtitle Token Query (TSTQ). The former aggregates global semantics from spatial contexts, while the latter queries subtitle-relevant tokens through temporal salience modeling. The output of S3 is projected into the language model’s embedding space and denoted as $\mathbf{F}^{out}$ after concatenation. This representation, conditioned on an instruction prompt $\mathcal{P}$, is subsequently fed into the large language model to produce the subtitle sequence $\mathcal{S}={S_{0},\ldots,S_{T-1}}$, where each subtitle $S_{i}$ is formatted as $\langle b_{i}\rangle\langle e_{i}\rangle s_{i}$”, with $b_{i}$ and $e_{i}$ denoting its start and end frames.
In the following subsections, we describe the vision encoder, the proposed S3 Module, and the large language model in detail.
III-B Vision Encoder
@@
In recent years, various Transformer-based vision backbones are proposed, including ViT [dosovitskiy2020image], Swin-Transformer [liu2021swin], EViT [liang2022not], Dynamic ViT [rao2021dynamicvit], etc. Among these Transformer-based backbones, we select Swin-Transformer pretrained in Donut [kim2021donut] as the vision encoder of EVE for two reasons: (1) Better initialization for OCR. Since the Swin-Transformer backbone of Donut has been pre-trained on numerous document datasets for multiple OCR tasks, it can provide better initialization to enhancing the text perception ability of EVE. In addition, based on the pre-training Swin-Transformer of Donut, our model can converge faster during the pre-training phase. (2) Higher time efficiency. Benefiting from the introduced shifted window-based self-attention module in Swin-Transformer, it shows better time efficiency, which can accelerate extracting visual features of frames in videos.
III-C Spatiotemporal Subtitle-Salient Module
The proposed S3 module comprises the following two branches:
III-C1 Spatial Semantic Context Aggregate
Subtitle generation is not a pure text-recognition problem but is inherently guided by the visual semantics of each frame. Movie and short-video subtitles often correspond to high-level scene cues—such as character interactions, camera composition, and background layout—that provide strong prior information about whether subtitles appear and how they should be phrased. To capture such global semantics in an efficient and robust manner, we introduce the Spatial Semantic Context Aggregate (SSCA) branch.
SSCA summarizes the frame feature $\mathbf{F}_{i}$ into a compact layout-aware representation $\mathbf{F}^{v}_{i}\in\mathbb{R}^{p\times p\times C}$ using spatial average pooling:
$\mathbf{F}^{v}_{i}=\texttt{AvgPool}(\mathbf{F}_{i}).$ (1)
Although the operation is simple, it serves a clear and task-aligned role: it suppresses low-level background noise while preserving high-level semantic structure—such as scene geometry, speaker regions, and global motion cues—that influence subtitle presence and content. SSCA thus provides a stable global prior that complements the text-focused modeling of the TSTQ branch.
III-C2 Temporal Subtitle Token Query
While SSCA captures holistic scene cues, accurate subtitle extraction further requires identifying fine-grained visual evidence related to on-screen text. To this end, we adopt a lightweight query-based design inspired by Q-Former [li2023blip]. A set of learnable query embeddings $\mathbf{Q}$ are used as latent probes that attend to subtitle-salient regions in the dense feature map. Each query specializes in capturing complementary cues—such as text contours, stroke patterns, or high-contrast regions—by performing cross-attention with $\mathbf{F}_{i}$:
$\mathbf{F}^{t}_{i}=\texttt{CrossAttention}(\mathbf{Q},\mathbf{F}_{i}).$ (2)
Unlike general vision–language alignment, subtitle extraction requires strong temporal stability because subtitles usually persist across several frames and gradually fade in or out. To enhance robustness, we extend the attention context to a temporal window consisting of the current frame and its neighbors:
$\mathbf{F}^{t}_{i}=\texttt{CrossAttention}(\mathbf{Q},[\mathbf{F}_{i-1},\mathbf{F}_{i},\mathbf{F}_{i+1}]).$ (3)
This temporal aggregation enables TSTQ to track slowly transitioning subtitle patterns, reduce frame-wise inconsistencies, and generate subtitle-related tokens that are both temporally consistent and visually discriminative.
III-C3 Feature Alignment and Fusion
The SSCA and TSTQ branches provide two complementary perspectives on subtitle perception: SSCA encodes global spatial semantics, while TSTQ extracts localized, temporally coherent subtitle tokens. Although both originate from the same visual encoder, their feature distributions vary substantially—one emphasizes holistic context, and the other emphasizes fine-grained text evidence.
To align these heterogeneous representations with LLM, we introduce the separate projector (SP) strategy :
$\hat{\mathbf{F}}^{v}_{i}=\texttt{Proj}_{v}(\mathbf{F}^{v}_{i}),\quad\hat{\mathbf{F}}^{t}_{i}=\texttt{Proj}_{t}(\mathbf{F}^{t}_{i}).$ (4)
The separate projector strategy prevents feature interference and preserves the modality-specific characteristics essential for subtitle reasoning.
Finally, the projected global semantics and subtitle-salient tokens across all frames are interleaved to form the unified output sequence:
$\mathbf{F}^{out}=[\hat{\mathbf{F}}^{v}_{0},\hat{\mathbf{F}}^{t}_{0},...,\hat{\mathbf{F}}^{v}_{N-1},\hat{\mathbf{F}}^{t}_{N-1}].$ (5)
This fused representation integrates both spatial and temporal cues in a subtitle-aware manner, enabling the language model to produce temporally aligned and semantically faithful subtitles.